OpenStudy (anonymous):

how do you solve this : -2(3x-4)+y=0

OpenStudy (anonymous):

step by step because i want to learn how to do it.

OpenStudy (anonymous):

Distribute the -2 into the ( ) first

OpenStudy (anonymous):

ok, so i got: -6x+4+y=0

OpenStudy (anonymous):

*+8

OpenStudy (anonymous):

yes Do you want to solve for y=mx+b

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Move the -6x and 4 to the other side of the equation.

OpenStudy (anonymous):

i thought it was 8

OpenStudy (anonymous):

because u had to distribute -2 to both 3x and -4

OpenStudy (anonymous):

sorry. I was looking at the wrong response..Move the -6x & 8 to the other side

OpenStudy (anonymous):

hold on, after distributing the -2, i got: (-6x+8)+y=0

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

so do i get: 8y=6x+0

OpenStudy (anonymous):

or is it wrong?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

oops -6x + 6x + 8 + y = 0+6x

OpenStudy (anonymous):

so 8+y=0+6x

OpenStudy (anonymous):

8 + y = 6x 8 - 8 + y = 6x - 8 y = 6x - 8

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

yes!thank u :)

OpenStudy (anonymous):

honestly, thank u very much.

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

how would u graph it thought?

OpenStudy (anonymous):